Find the domain of the function. (Enter your answer using interval notation.)
f(x) =
5x^3 − 5/ x^2 + 2x − 15

Answer by lwsshak3(11628)

Find the domain of the function. (Enter your answer using interval notation.)
f(x) = 5x^3 − 5/ x^2 + 2x − 15
function is undefined if denominator=0
so, set denominator=0, and solve for x.
x^2+2x−15=0
(x+5)(x-3)=0
x=-5
or
x=3
domain: (-∞,-5) U (-5,3) U (3,∞)